40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13870     Entry time: Sun May 20 23:43:50 2018     In reply to: 13770     Reply to this: 13878
Author: gautam 
Type: Update 
Category: IOO 
Subject: Coil driver noise re-measurement 

Summary:

In the IMC actuation chain, it looks like the MC1/MC3 de-whitening boards, and also all three MC optics' coil driver boards, are showing higher noise than expected from LISO modeling. One possible candidate is thick film resistors on the coil driver boards. The plan is to debug these further by pulling the board out of the Eurocrate and investigating on the electronics bench.

Why bother? Mainly because I want to see how good the IR ALS noise is, and currently, the PSL frequency noise is causing the measurement to be worse than references taken from previous known good times.

Details:

Sometime ago, rana suggested to me that I should do this measurement more systematically.

  • Attachments #1, #2 and #3 show noise measurements in various conditions for MC1, MC2 and MC3 respectively.
  • In the above three attachments, I stitched together multiple spans from the SR785, and so the bin-width is different. The data is downloaded from the analyzer normalized by the bin-width (PSD units).
  • The roll-off at ~800Hz in the orange trace for MC1 and MC3 is consistent with an 800 Hz LPF that was used for anti-image filtering from the old 2 kHz era.
  • While it may look like the analog de-whitening isn't being switched on in some of these plots, I confirmed by transfer function measurement that it is indeed switching.
  • Data used to make these plots are in Attachment #4. Unfortunately, the code requires some of my personal plotting librariesno and so I'm not uploading it.
  • Sketch of measurement setup is shown in Attachment #5. It is not indicated in the schematic, but the SR560 was operated in battery mode for this measurement.
  • For MC1, I did the additional measurement of terminating the LEMO input to the coil driver and measuring (what should have been) just the coil driver electronics noise. But this measurement doesn't look very clean, and hence, the decision to pull the board out to continue debugging.
  • While at 1X6, Rana tightened the LEMO connectors going to MC1. We should opportunistically do MC2 and MC3 as well.
  • Some changes to be made:
    • Thick film ---> thin film.
    • Reroute HPF-ed back-plane Vmon output to the front panel LEMO.

I've now restored all the wiring at 1X6 to their state before this work.

Attachment 1: MC1_coilDriver.pdf  512 kB  Uploaded Mon May 21 00:51:19 2018  | Hide | Hide all
MC1_coilDriver.pdf
Attachment 2: MC2_coilDriver.pdf  443 kB  Uploaded Mon May 21 00:51:33 2018  | Hide | Hide all
MC2_coilDriver.pdf
Attachment 3: MC3_coilDriver.pdf  444 kB  Uploaded Mon May 21 00:51:50 2018  | Hide | Hide all
MC3_coilDriver.pdf
Attachment 4: MC_coilDriverNoises.tgz  695 kB  Uploaded Mon May 21 01:01:23 2018
Attachment 5: ActuationChainNoiseMeas.pdf  30 kB  Uploaded Mon May 21 01:56:08 2018  | Hide | Hide all
ActuationChainNoiseMeas.pdf
ELOG V3.1.3-